# Self-Correcting Decoding with Generative Feedback for Mitigating Hallucinations in Large Vision-Language Models

### v1

@@ 4.1 Experimental Settings
| MS-COCO | Random | Regular | 83.13 | 81.94 | 83.44 | 83.07 | 83.02 | 83.08 | 87.43 | 93.56 | 86.48 |
@@
| A-OKVQA | Random | Regular | 81.90 | 76.63 | 83.53 | 80.63 | 76.82 | 81.92 | 86.27 | 90.66 | 85.48 |
@@
Baselines. As a simple baseline, we include results from regular decoding, where the next token is sampled directly from the post-softmax probability distribution. Additionally, we compare the performance of our method with three state-of-the-art decoding approaches: VCD (Leng et al., 2024), M3ID (Favero et al., 2024), and RITUAL (Woo et al., 2024). For evaluations on the CHAIR (Rohrbach et al., 2018) and MME-Hallucination (Fu et al., 2023) benchmark, we further include comparisons with Woodpecker (Chen et al., 2024d), HALC (Chen et al., 2024d), DoLa (Chuang et al., 2024) and OPERA (Huang et al., 2024). We report the performance of these baselines based on our re-implementation using their released code bases.
@@ 4.2 Results and Discussions
Results on POPE. In Table 1, we compare the performance of our method against other baselines on the POPE benchmark under three different negative sampling settings, across three datasets. As shown in the table, our method consistently outperforms other decoding methods on both LVLMs, achieving state-of-the-art accuracies across all 18 settings, with improvements of up to 5.24% in accuracy, 6.33% in precision, and 2.79% in F1 score compared to the second-best approach. This suggests that incorporating a generative reference enables the LVLMs to perceive more fine-grained visual details, thereby effectively addressing object hallucinations. Moreover, while most decoding methods tend to be overconfident in their responses, the self-correcting decoding mechanism in our method makes it more conservative in responding Yes, as evidenced by significantly higher precision across all settings. This highlights its enhanced performance in filtering out false positives and suppressing misinformation.
@@ Acknowledgements
This work has been funded in part by the Army Research Laboratory (ARL) award W911NF-23-2-0007, DARPA award FA8750-23-2-1015, and ONR award N00014-23-1-2840. MM and LPM are partially supported by Meta and National Institutes of Health awards R01MH125740, R01MH132225, and R21MH130767. RS is supported in part by the ONR grant N00014-23-1-2368.
@@
Our work focuses on developing methods to mitigate hallucinations in large vision-language models, aiming to enhance the reliability of AI-generated content. Our research does not involve human subjects, sensitive data, or any practices that pose privacy or security concerns. Additionally, we discuss the broader ethical and societal implications of this work in Section A of the Appendix.
@@ C.8 Scaling Up the LVLMs
| MS-COCO | Random | Regular | 82.53 | 78.57 | 89.47 | 83.67 |

### v2

@@ 4 Experiments
MS-COCO |
Random | Regular | 83.13 | 81.94 | 83.44 | 83.07 | 83.02 | 83.08 | 87.43 | 93.56 | 86.48 |
@@
A-OKVQA |
Random | Regular | 81.90 | 76.63 | 83.53 | 80.63 | 76.82 | 81.92 | 86.27 | 90.66 | 85.48 |
@@
Baselines. As a simple baseline, we include results from regular decoding, where the next token is sampled directly from the post-softmax probability distribution. Additionally, we compare the performance of our method with three state-of-the-art decoding approaches: VCD (Leng et al., 2024), M3ID (Favero et al., 2024), and RITUAL (Woo et al., 2024). For evaluations on the CHAIR (Rohrbach et al., 2018) and MME-Hallucination (Fu et al., 2023) benchmark, we further include comparisons with Woodpecker (Yin et al., 2023), HALC (Chen et al., 2024d), DoLa (Chuang et al., 2024) and OPERA (Huang et al., 2024). We report the performance of these baselines based on our re-implementation using their released code bases.
@@ 4.1 Results and Discussions
Results on POPE. In Table 1, we compare the performance of our method against other baselines on the POPE benchmark under three different negative sampling settings, across three datasets. As shown in the table, our method consistently outperforms other decoding methods on three LVLMs, achieving state-of-the-art accuracies across all settings, with improvements of up to 5.24% in accuracy, 6.33% in precision, and 2.79% in F1 score compared to the second-best approach. This suggests that incorporating a generative reference enables the LVLMs to perceive more fine-grained visual details, thereby effectively addressing object hallucinations. Moreover, while most decoding methods tend to be overconfident in their responses, the self-correcting decoding mechanism in our method makes it more conservative in responding Yes, as evidenced by significantly higher precision across all settings. This highlights its enhanced performance in filtering out false positives and suppressing misinformation.
@@ Acknowledgements
This work has been funded in part by the Army Research Laboratory (ARL) award W911NF-23-2-0007 and W911QX-24-F-0049, DARPA award FA8750-23-2-1015, and ONR award N00014-23-1-2840. MM and LPM are partially supported by Meta and National Institutes of Health awards R01MH125740, R01MH132225, and R21MH130767. RS is supported in part by the ONR grant N00014-23-1-2368.
@@
Our work focuses on enhancing the reliability of current large vision-language models. Our research does not involve human subjects, sensitive data, or any practices that pose privacy or security concerns. We also discuss the broader ethical and societal implications of this work in Appendix A.
@@ C.8 Scaling Up the LVLMs
MS-COCO |
Random | Regular | 82.53 | 78.57 | 89.47 | 83.67 |